If f and g are arbitrary function of their respective arguments, show that u= f( x -vt+i alpha y ) is a solution of (d^2u/dx^2 ) + (d^2u/dy^2) = (1/c^2) (d^2u/dt^2), where (alpha)^2 = 1-(v^2/c^2)?
1
Expert's answer
2019-09-27T09:04:56-0400

Consider "u=f(x-vt+i\alpha y)."

Differentiating "u" partially with respect to "x," we get


"{\\partial\tu \\over \\partial x}=f'(x-vt+i\\alpha y)""{\\partial^2 \tu \\over \\partial x^2 }=f''(x-vt+i\\alpha y)"

Differentiating "u" partially with respect to "y," we get


"{\\partial\tu \\over \\partial y}=f'(x-vt+i\\alpha y)\\cdot i\\alpha""{\\partial^2\tu \\over \\partial y^2 }=f''(x-vt+i\\alpha y)\\cdot (i\\alpha)^2"

Then


"{\\partial^2\tu \\over \\partial y^2 }=-\\alpha^2{\\partial^2 u \\over \\partial x^2 }"

Differentiating "u" partially with respect to "t," we get


"{\\partial\tu \\over \\partial t}=f'(x-vt+i\\alpha y)\\cdot (-v)""{\partial^2u \over \partial t^2 }=f''(x-vt+i\alpha y)\cdot (-v)^2"

Then


"{\partial^2u \over \partial t^2 }=v^2 {\partial^2u \over \partial x^2 }"

Substitute into the given equation


"{\partial^2u \over \partial x^2 }-\alpha^2{\partial^2u \over \partial x^2 }={v^2 \over c^2}\cdot {\partial^2u \over \partial x^2 }"

Take

"\alpha^2=1-{v^2 \over c^2}"

"{\partial^2u \over \partial x^2 }-\bigg(1-{v^2 \over c^2}\bigg){\partial^2u \over \partial x^2 }={v^2 \over c^2}\cdot {\partial^2u \over \partial x^2 }"

"{v^2 \over c^2}\cdot {\partial^2u \over \partial x^2 }={v^2 \over c^2}\cdot {\partial^2u \over \partial x^2 }"

"{\partial^2u \over \partial x^2 }= {\partial^2u \over \partial x^2 },\ True"

Therefore, "u=f(x-vt+i\alpha y)" satisfies the given differential equation and hence the solution of the equation.